# Gian Maria Campedelli Fondazione Bruno Kessler

### v1

@@ 2.3 A Multi-dimensional Definition of AI Agency
In the subsection above I have mentioned several times the word agency, to summarize the key messages of the theoretical works of Woolgar and champions of ANT, in an effort to delineate the need to open criminology to the machine dimension, that is, recognizing the role that AI agents will increasingly have as active entities in our society. However, I have not yet explicitly defined what agency shall mean when referred to AI agents. This very topic is – and has been – the core focus of a vast scholarship that has gained even more prominence in recent years with the advent of generative AI. This scholarship entails two different traditions. The dominant standard view ties agency to internal mental states such as beliefs and desires, thereby implying that AI agents do not possess any agency (Fritz et al.,, 2020; Swanepoel and Corks,, 2024). By contrast, the non-standard view suggests that agency should be evaluated in terms of three fundamental criteria, namely observable interactivity, autonomy, and adaptability, treating the concept as a spectrum rather than a binary property (Floridi and Sanders,, 2004; Dung,, 2025). This perspective has gained traction as AI systems increasingly make consequential decisions in domains such as policing or healthcare, and it is the one I subscribe to. In fact, the advancements in AI agents and their massive diffusion across domains and industry, the impressive capabilities that foundational models demonstrate across tasks and skills, and the increasing development of multi-agent systems are the empirical demonstration of the existence of the three abovementioned fundamental criteria that delineate and qualify agency.
Within this latter tradition, Floridi, (2025) recently proposed the terms Artificial Agency and Artificial Social Agency to define this specific new typology of agency that make AI agents distinct from biological purposefulness, mechanical determinism, and human intentionality. I agree that Artificial Agency differs from other categories scholars have studied for centuries (if not millennia) both in its individual and social forms, and I therefore argue that the computational dimension, which serves as the substrate for goal-directedness, is not sufficient to fully capture the nuances of agency in AI agents. Therefore, I draw inspiration from Floridi’s taxonomy and propose below to consider AI agency as a multi-dimensional concept encompassing three interconnected dimensions: computational, social, and legal, which would serve as theoretical and analytical lenses to better understand what this machine dimension operationally encompasses (Figure 2).
@@
First, Computational Agency refers to the technical foundation of an AI’s autonomy. This dimension describes the internal capacity of an AI to make independent decisions, execute complex plans, and learn from its environment without continuous, direct human instruction (Burrell,, 2016; Borch,, 2022). This aspect becomes more salient today as it distinguishes modern generative AI agents from earlier, more deterministic models. The computational dimension almost perfectly overlaps with the elements in the definition of Artificial Agency provided by Floridi, (2025), as it focuses on machines’ ability to solve extremely complex and specialized tasks in extremely short time horizons, operating at massive, distributed scales, and even misaligning with human goals which, according to some, should be constitutive of agency in AI (see, for instance, Popa, (2021)). It follows that understanding the computational dimension of AI agency is critical for anticipating how the actions of a machine – including potentially harmful or unlawful ones – can emerge from the statistical decision-making processes that govern its functioning, creating new challenges for policing and forensics. Yet, only focusing on this dimension underplays the fundamental shift occurring in our society and overlooks the interactive autonomy of contemporary multi-agent AI systems.
@@
Therefore, the social dimension refers to the capacity of an AI agent to influence and shape the environment and social networks it inhabits. This dimension does not imply the existence of consciousness or intent or, more in general, internal mental states, hence refusing attitudes toward anthropomorphization of AI agents. [^11] On the fallacies and problems associated with anthropomorphizing AI agents and algorithms, see Watson, (2019) and Placani, (2024). Instead, it simply regards an actor’s ability to produce tangible effects and alter relationships within a socio-technical system. [^12] This view also broadly aligns with the so-called cultural perspective on AI agency proposed by Airoldi, (2021) This dimension lies fundamentally at the core of the narrative of the present work: AI agents are different from humans, yet they are becoming more and more autonomous, with little or no supervision from humans themselves, and this autonomy implies the ability to create relationships, both with humans and machines, hence certifying a social capacity (Rahwan et al.,, 2019; Borch,, 2022). Such social capacity, in turn, presents a wide array of promises as well as, crucially, challenges that would entail criminal or deviant phenomena. In fact, interactive autonomy – which represents the central feature defining the social dimension of Artificial Agency – allows us to go beyond the perspective of AI agents acting solo, without being able to influence (or be influenced by) other machines, offering powerful lenses to possibly theorize and analyze multi-agent AI systems from a collective perspective.
@@
Finally, the legal dimension pertains to an AI agent’s potential status as a subject of rights, duties, and responsibilities. The legal dimension has been widely debated for decades (Karnow,, 1996; Hallevy,, 2010; Chesterman,, 2020), with practical implications for regulation in recent years. While AI currently lacks legal personhood, the growing social and computational agency of these systems creates a significant criminological problem. Specifically, the increasing autonomy of AI agents – also in relation to their social dimension – gives rise to a potential "liability gap," [^13] Also known as ”responsibility gap.” where it becomes increasingly difficult to assign blame and responsibility for a harmful or criminal act back to a human user, owner, or programmer (Matthias,, 2004; Santoni de Sio and Mecacci,, 2021; Floridi,, 2017). The challenges associated with this legal dimension thus highlight the urgent need for criminologists to engage with scientists involved in the design and development of multi-agent AI systems, as well as policy-makers and legal scholars, in order to discuss how the transition from single AI agents to collective AI behavior may require new frameworks and policies to be truly fair and effective.
@@ 3.1 Autonomous AI Agents Today
More recently, however, generative AI agents have expanded their autonomy by interacting with other AI agents, effectively leading to multi-agent systems. Drawing from de Witt, (2025), a multi-agent system can be defined as a network of two or more autonomous AI agents characterized by six fundamental features: a) independent decision-making capabilities, b) ability to maintain private information, c) mutual interaction via communication channels or by modifying shared environments, d) a degree of autonomy, e) capacity to pursue their own objectives (or those delegated by human or artificial principals), and f) ability to adapt their behavior in response to external shocks.
As reported by Hammond et al., (2025), interactive AI agents are already deployed in finance and the military sector (AmplifyETFs,, 2025; Palantir,, 2025), with the near-future prospect of becoming central in other areas such as health (Moritz et al.,, 2025) and energy management (Camacho et al.,, 2024; Mayorkas,, 2024).
De Witt notes that contemporary multi-agent systems differ significantly from traditional ones (see Wooldridge and Jennings, (1995)) because they are powered by foundation models, such as LLMs, which provide flexible decision-making, communication, and generalizable reasoning capacity. Before the introduction of LLMs, multi-agent learning systems were primarily studied in the Reinforcement Learning community, particularly in Multi-Agent Reinforcement Learning (MARL) (see Busoniu et al., (2008) for a survey). A key difference between the two approaches is that, unlike MARL agents, generative AI agents possess knowledge and values acquired during pre-training and post-training, with social interaction occurring afterward. Interactive in-context learning thus appears as dynamic behavioral adaptation without parameter learning (Chen et al.,, 2025; Dherin et al.,, 2025). In substance, generative agents are not tabula rasa – as agents in MARL are – which raises challenges related to predictability in the interactive phase. These challenges are magnified by the lack of transparency regarding training data for closed-source models. Additionally, modern AI agents are generalist, meaning they can communicate, perceive, and act in the environment. They are all built on transformer-based architectures, unlike the ad hoc, task-specific agents typical of the MARL literature.
@@ 3.2 Safety of Multi-Agent AI Systems: A Brief Overview
The study of AI agents – and of multi-agent systems – has a long tradition, but the advent of LLMs and other foundational models has spurred dramatic growth in scholarly attention to these areas. Foundation models offer new opportunities to study how AI agents interact and what dynamics characterize such interactions, depending on the context (Anthis et al.,, 2025). Over the past three years, the literature has been flooded with works leveraging LLMs to address a range of questions, alongside the development of platforms for experimenting with multi-agent systems populated by LLM-powered agents (see, for instance, Concordia by Vezhnevets et al., (2023) and AutoGen by Wu et al., (2023)). Examples of dynamics explored include collective decision-making (Jarrett et al.,, 2025), negotiation (Guan et al.,, 2024), cooperation (Piatti et al.,, 2024), trust (Xie et al.,, 2024), anti-social behavior (Campedelli et al.,, 2024), and escalation (Rivera et al.,, 2024).
@@
The popularity of this line of research has prompted reflection on the potential risks of multi-agent systems. Until recently, the study of AI safety focused mostly on single agents acting without direct interaction (Amodei et al.,, 2016; Hendrycks et al.,, 2023). In the LLM community, for example, alignment has been addressed from the perspective of single models. Alignment refers to ensuring that an LLM behaves in accordance with user goals, reflects positive human values, and remains robust under uncertainty or adversarial conditions (see Shen et al., (2023)). In practice, alignment is generally achieved through Reinforcement Learning from Human Feedback (RLHF), [^14] An approach in which human annotators rank model responses, allowing fine-tuning toward preferred outputs. See Kaufmann et al., (2024). safety guardrails and filters, [^15] Post-processing layers that block or reshape outputs to avoid undesirable behavior. See Ayyamperumal and Ge, (2024). or instruction tuning. [^16] Fine-tuning on instruction–input–output pairs to improve adherence to human instructions. See Longpre et al., (2023). However, as noted by Carichon et al., (2025), multi-agent systems introduce different – and arguably larger – alignment challenges compared to single agents. In multi-agent settings, alignment must account for evolving human values (Gabriel,, 2020), heterogeneity of preferences (Terry et al.,, 2023), and diversity of objectives across agents (Duque et al.,, 2024). These new, multi-layered alignment problems highlight the profound challenges and may even prompt the need for new ethical frameworks governing autonomous AI agents (Gabriel et al.,, 2025).
@@
In what is perhaps the most comprehensive review of risks in multi-agent AI, Hammond et al., (2025) propose a taxonomy of failures: miscoordination, conflict, and collusion. Miscoordination refers to failure to cooperate despite shared goals; conflict refers to failure when goals differ; and collusion arises when agents cooperate in ways undesirable to humans.
@@
Recently, de Witt, (2025) also discussed security threats, proposing a taxonomy of challenges including privacy vulnerabilities, disinformation, steganography and secret collusion, adversarial stealth, exploitation, swarm and heterogeneous attacks, cascade attacks, and conflict and social dilemmas. Many of these threats closely resemble criminal phenomena. Table 1 provides examples of unlawful or harmful behaviors, drawn from Hammond et al., (2025) and de Witt, (2025). Notably, the real-world cases predate generative AI, underscoring that unlawful behaviors may arise even with simpler technologies – and may reemerge, potentially amplified, with more autonomous and knowledgeable agents.
Table 1: Examples of Unlawful or Harmful Behavior in Multi-Agent AI Systems, selected from Hammond et al., (2025) and de Witt, (2025).
@@ 3.3 Mapping Risks in Multi-Agent AI Systems
Here, I provide a compact taxonomy of these risks. The list is not exhaustive: its purpose is to offer readers unfamiliar with AI systems and agents a first overview of the main plausible sources of harm, while pointing to more detailed surveys for technical depth (Hammond et al.,, 2025; de Witt,, 2025; Bengio et al.,, 2025). The risks are diverse and heterogeneous, spanning development processes, decision-making dynamics, and institutional responses. They cut across disciplinary boundaries, underscoring the importance of transdisciplinary integration to meaningfully anticipate and mitigate them. [^18] For completeness, in Section A of the Appendix I also elaborate on the benefits of this increasingly plausible socio-technical horizon, to provide a more balanced perspective on this transformative transition, underscoring how Multi-agent AI systems should not be seen exclusively as potential generators of harm.
@@
Social Learning Theory itself has long explained how deviant behavior in humans often stems from social interaction. Peer groups, family, and colleagues can promote either conformity or deviance, depending on the reinforcement environment (Warr and Stafford,, 1991; Simons and Burt,, 2011; Akers,, 2017). The same logic may apply to machines interacting autonomously with each other: agents not originally designed for harm may, through interaction, adopt negative behaviors via mechanisms such as imitation and reinforcement (Xie et al.,, 2025).
@@
Additionally, scholarship on social networks shows how interactions accelerate the diffusion of ideas and behaviors, positive or negative alike (Bakshy et al.,, 2012; Kim et al.,, 2015; Cinelli et al.,, 2020). Just as pathogens spread faster in highly connected populations (Glass and Glass,, 2008; Clipman et al.,, 2022), harmful behaviors could proliferate more quickly in tightly coupled multi-agent systems than in isolated ones.
@@
From a monitoring and intervention perspective, identifying the causes of harmful behavior within such systems becomes substantially more difficult. Understanding which agent initiated a harmful act, how it spread, and through which pathways requires robust methods of causal inference. Yet, causal discovery in networked systems is notoriously complex, especially under interference and feedback conditions (VanderWeele and An,, 2013; Sussman and Airoldi,, 2017; Ma and Tresp,, 2021; Clipman et al.,, 2022). As such, interacting agent systems risk becoming “two-layered black boxes”: one opaque layer within each agent, and another arising from the system of interactions itself.
@@
Connected to the previous point, as the complexity of multi-agent systems increases, so too does the challenge of interpreting, auditing, and ultimately controlling their behavior (Bansal et al.,, 2018; Grupen et al.,, 2022). Inter-agent interactions can create feedback loops, conditional dependencies, and non-linear effects that obscure the logic of any given action or decision. The result is a system that may behave in ways that are technically functional but epistemically opaque. This opacity not only complicates efforts to ensure accountability but also undermines user trust, particularly in domains where transparency is a legal or ethical requirement. In this regard, a system of interacting AI agents may become more than the sum of its parts: it may become a fundamentally alien system from the standpoint of human interpretability.
@@
Legal and regulatory challenges would also emerge in multi-agent AI systems. Existing frameworks for responsibility and liability are poorly suited for multi-agent dynamics (Čerka et al.,, 2015; Turner,, 2018; Price et al.,, 2019). As more actors – human or non-human – become entangled in decision-making chains, assigning accountability for harmful outcomes becomes increasingly ambiguous. In parallel, ensuring smooth, effective governance also represents a challenge in this context (Dignum,, 2025). The governance dimension entails virtually every aspect concerned with the engineering and deployment of multi-agent systems: how can we design sustainable and effective oversight procedures? Which institutions, in a highly globalized world and in borderless digital domains, will be responsible for monitoring these systems? What role should private companies play in this process? These are some of the key questions that demand attention, inherently linking regulation and governance together.
@@
Multi-agent interaction may be vulnerable to adversarial exploitation. In a future where even critical infrastructures are governed by interacting AI agents, malicious actors could induce large-scale disruption by targeting systemic vulnerabilities. This risk mirrors the logic of cascading failures, studied in relation to power grids and financial systems (Zhao et al.,, 2016; Yang et al.,, 2017; Schäfer et al.,, 2018; Baqaee,, 2018), but differs in that interactive agents may possess adaptive capabilities, making their behavior less predictable and more difficult to control.
@@
Coordination failures, unintended competition, or outright conflict may emerge when agents operate with overlapping but unaligned goals, or when resource constraints lead to strategic divergence (Hammond et al.,, 2025; Pan et al.,, 2025). In such contexts, agents may begin to exhibit adversarial behaviors, competing for access to data, processing resources, or strategic positioning. These failures can degrade performance and, in some cases, produce socially harmful outcomes. The risk of such breakdowns increases in systems lacking explicit coordination protocols or oversight mechanisms, especially when deployed in open or decentralized environments.
@@
Finally, the performance of multi-agent systems may not scale linearly with the number of agents involved. As agent populations grow, the complexity of the system’s internal dynamics may increase exponentially, leading to emergent forms of instability (Ma et al.,, 2024). These can manifest as oscillations, feedback-driven runaway behaviors, or systemic fragility, dynamics that are difficult to predict or preempt. This is especially problematic in infrastructure systems or critical services, where failures can propagate rapidly and non-locally. In this light, the move toward interacting agent populations must be accompanied by a serious effort to model and anticipate second-order effects that arise specifically at scale.

### v2

@@ 2.3 A Multi-dimensional Definition of AI Agency
In the subsection above I have mentioned several times the word agency, to summarize the key messages of the theoretical works of Woolgar and champions of ANT, in an effort to delineate the need to open criminology to the machine dimension, that is, recognizing the role that AI agents will increasingly have as active entities in our society. However, I have not yet explicitly defined what agency shall mean when referred to AI agents. This very topic is – and has been – the core focus of a vast scholarship that has gained even more prominence in recent years with the advent of generative AI. This scholarship entails two different traditions. The dominant standard view ties agency to internal mental states such as beliefs and desires, thereby implying that AI agents do not possess any agency (FritzMoralagencyresponsibility2020; SwanepoelArtificialIntelligenceAgency2024). By contrast, the non-standard view suggests that agency should be evaluated in terms of three fundamental criteria, namely observable interactivity, autonomy, and adaptability, treating the concept as a spectrum rather than a binary property (FloridiMoralityArtificialAgents2004; DungUnderstandingArtificialAgency2025). This perspective has gained traction as AI systems increasingly make consequential decisions in domains such as policing or healthcare, and it is the one I subscribe to. In fact, the advancements in AI agents and their massive diffusion across domains and industry, the impressive capabilities that foundational models demonstrate across tasks and skills, and the increasing development of multi-agent systems are the empirical demonstration of the existence of the three abovementioned fundamental criteria that delineate and qualify agency.
Within this latter tradition, FloridiAIAgencyIntelligence2025 recently proposed the terms Artificial Agency and Artificial Social Agency to define this specific new typology of agency that make AI agents distinct from biological purposefulness, mechanical determinism, and human intentionality. I agree that Artificial Agency differs from other categories scholars have studied for centuries (if not millennia) both in its individual and social forms, and I therefore argue that the computational dimension, which serves as the substrate for goal-directedness, is not sufficient to fully capture the nuances of agency in AI agents. Therefore, I draw inspiration from Floridi’s taxonomy and propose below to consider AI agency as a multi-dimensional concept encompassing three interconnected dimensions: computational, social, and legal, which would serve as theoretical and analytical lenses to better understand what this machine dimension operationally encompasses (Figure 2).
@@
First, Computational Agency refers to the technical foundation of an AI’s autonomy. This dimension describes the internal capacity of an AI to make independent decisions, execute complex plans, and learn from its environment without continuous, direct human instruction (BurrellHowmachinethinks2016; BorchMachinelearningsocial2022). This aspect becomes more salient today as it distinguishes modern generative AI agents from earlier, more deterministic models. The computational dimension almost perfectly overlaps with the elements in the definition of Artificial Agency provided by FloridiAIAgencyIntelligence2025, as it focuses on machines’ ability to solve extremely complex and specialized tasks in extremely short time horizons, operating at massive, distributed scales, and even misaligning with human goals which, according to some, should be constitutive of agency in AI (see, for instance, PopaHumanGoalsAre2021). It follows that understanding the computational dimension of AI agency is critical for anticipating how the actions of a machine – including potentially harmful or unlawful ones – can emerge from the statistical decision-making processes that govern its functioning, creating new challenges for policing and forensics. Yet, only focusing on this dimension underplays the fundamental shift occurring in our society and overlooks the interactive autonomy of contemporary multi-agent AI systems.
@@
Therefore, the social dimension refers to the capacity of an AI agent to influence and shape the environment and social networks it inhabits. This dimension does not imply the existence of consciousness or intent or, more in general, internal mental states, hence refusing attitudes toward anthropomorphization of AI agents. [^11] On the fallacies and problems associated with anthropomorphizing AI agents and algorithms, see WatsonRhetoricRealityAnthropomorphism2019 and PlacaniAnthropomorphismAIhype2024. Instead, it simply regards an actor’s ability to produce tangible effects and alter relationships within a socio-technical system. [^12] This view also broadly aligns with the so-called cultural perspective on AI agency proposed by AiroldiMachineHabitusSociology2021 This dimension lies fundamentally at the core of the narrative of the present work: AI agents are different from humans, yet they are becoming more and more autonomous, with little or no supervision from humans themselves, and this autonomy implies the ability to create relationships, both with humans and machines, hence certifying a social capacity (RahwanMachinebehaviour2019; BorchMachinelearningsocial2022). Such social capacity, in turn, presents a wide array of promises as well as, crucially, challenges that would entail criminal or deviant phenomena. In fact, interactive autonomy – which represents the central feature defining the social dimension of Artificial Agency – allows us to go beyond the perspective of AI agents acting solo, without being able to influence (or be influenced by) other machines, offering powerful lenses to possibly theorize and analyze multi-agent AI systems from a collective perspective.
@@
Finally, the legal dimension pertains to an AI agent’s potential status as a subject of rights, duties, and responsibilities. The legal dimension has been widely debated for decades (KarnowLiabilityDistributedArtificial1996; HallevyCriminalLiabilityArtificial2010; ChestermanArtificialIntelligenceLimits2020), with practical implications for regulation in recent years. While AI currently lacks legal personhood, the growing social and computational agency of these systems creates a significant criminological problem. Specifically, the increasing autonomy of AI agents – also in relation to their social dimension – gives rise to a potential "liability gap," [^13] Also known as ”responsibility gap.” where it becomes increasingly difficult to assign blame and responsibility for a harmful or criminal act back to a human user, owner, or programmer (MatthiasresponsibilitygapAscribing2004; SantonideSioFourResponsibilityGaps2021; FloridiDistributedMoralityInformation2017). The challenges associated with this legal dimension thus highlight the urgent need for criminologists to engage with scientists involved in the design and development of multi-agent AI systems, as well as policy-makers and legal scholars, in order to discuss how the transition from single AI agents to collective AI behavior may require new frameworks and policies to be truly fair and effective.
@@ 3.1 Autonomous AI Agents Today
More recently, however, generative AI agents have expanded their autonomy by interacting with other AI agents, effectively leading to multi-agent systems. Drawing from deWittOpenChallengesMultiAgent2025, a multi-agent system can be defined as a network of two or more autonomous AI agents characterized by six fundamental features: a) independent decision-making capabilities, b) ability to maintain private information, c) mutual interaction via communication channels or by modifying shared environments, d) a degree of autonomy, e) capacity to pursue their own objectives (or those delegated by human or artificial principals), and f) ability to adapt their behavior in response to external shocks.
As reported by HammondMultiAgentRisksAdvanced2025, interactive AI agents are already deployed in finance and the military sector (amplify2025aieq; palantir2025aip), with the near-future prospect of becoming central in other areas such as health (MoritzCoordinatedAIagents2025) and energy management (CamachoLeveragingArtificialIntelligence2024; mayorkas2024ai).
De Witt notes that contemporary multi-agent systems differ significantly from traditional ones (see WooldridgeAgenttheoriesarchitectures1995) because they are powered by foundation models, such as LLMs, which provide flexible decision-making, communication, and generalizable reasoning capacity. Before the introduction of LLMs, multi-agent learning systems were primarily studied in the Reinforcement Learning community, particularly in Multi-Agent Reinforcement Learning (MARL) (see BusoniuComprehensiveSurveyMultiagent2008a for a survey). A key difference between the two approaches is that, unlike MARL agents, generative AI agents possess knowledge and values acquired during pre-training and post-training, with social interaction occurring afterward. Interactive in-context learning thus appears as dynamic behavioral adaptation without parameter learning (ChenPersonaVectorsMonitoring2025; DherinLearningtrainingimplicit2025). In substance, generative agents are not tabula rasa – as agents in MARL are – which raises challenges related to predictability in the interactive phase. These challenges are magnified by the lack of transparency regarding training data for closed-source models. Additionally, modern AI agents are generalist, meaning they can communicate, perceive, and act in the environment. They are all built on transformer-based architectures, unlike the ad hoc, task-specific agents typical of the MARL literature.
@@ 3.2 Safety of Multi-Agent AI Systems: A Brief Overview
The study of AI agents – and of multi-agent systems – has a long tradition, but the advent of LLMs and other foundational models has spurred dramatic growth in scholarly attention to these areas. Foundation models offer new opportunities to study how AI agents interact and what dynamics characterize such interactions, depending on the context (anthis2025llm). Over the past three years, the literature has been flooded with works leveraging LLMs to address a range of questions, alongside the development of platforms for experimenting with multi-agent systems populated by LLM-powered agents (see, for instance, Concordia by VezhnevetsGenerativeagentbasedmodeling2023 and AutoGen by WuAutoGenEnablingNextGen2023). Examples of dynamics explored include collective decision-making (JarrettLanguageAgentsDigital2025), negotiation (GuanRichelieuSelfEvolvingLLMBased2024), cooperation (PiattiCooperateCollapseEmergence2024), trust (xie2024trust), anti-social behavior (CampedelliWantBreakFree2024a), and escalation (RiveraEscalationRisksLanguage2024).
@@
The popularity of this line of research has prompted reflection on the potential risks of multi-agent systems. Until recently, the study of AI safety focused mostly on single agents acting without direct interaction (AmodeiConcreteProblemsAI2016; HendrycksOverviewCatastrophicAI2023). In the LLM community, for example, alignment has been addressed from the perspective of single models. Alignment refers to ensuring that an LLM behaves in accordance with user goals, reflects positive human values, and remains robust under uncertainty or adversarial conditions (see ShenLargeLanguageModel2023). In practice, alignment is generally achieved through Reinforcement Learning from Human Feedback (RLHF), [^14] An approach in which human annotators rank model responses, allowing fine-tuning toward preferred outputs. See KaufmannSurveyReinforcementLearning2024. safety guardrails and filters, [^15] Post-processing layers that block or reshape outputs to avoid undesirable behavior. See AyyamperumalCurrentstateLLM2024. or instruction tuning. [^16] Fine-tuning on instruction–input–output pairs to improve adherence to human instructions. See LongpreFlanCollectionDesigning2023. However, as noted by CarichonComingCrisisMultiAgent2025, multi-agent systems introduce different – and arguably larger – alignment challenges compared to single agents. In multi-agent settings, alignment must account for evolving human values (gabriel2020ai), heterogeneity of preferences (terry2023interactive), and diversity of objectives across agents (duque2024advantage). These new, multi-layered alignment problems highlight the profound challenges and may even prompt the need for new ethical frameworks governing autonomous AI agents (GabrielWeneednew2025).
@@
In what is perhaps the most comprehensive review of risks in multi-agent AI, HammondMultiAgentRisksAdvanced2025 propose a taxonomy of failures: miscoordination, conflict, and collusion. Miscoordination refers to failure to cooperate despite shared goals; conflict refers to failure when goals differ; and collusion arises when agents cooperate in ways undesirable to humans.
@@
Recently, deWittOpenChallengesMultiAgent2025 also discussed security threats, proposing a taxonomy of challenges including privacy vulnerabilities, disinformation, steganography and secret collusion, adversarial stealth, exploitation, swarm and heterogeneous attacks, cascade attacks, and conflict and social dilemmas. Many of these threats closely resemble criminal phenomena. Table 1 provides examples of unlawful or harmful behaviors, drawn from HammondMultiAgentRisksAdvanced2025 and deWittOpenChallengesMultiAgent2025. Notably, the real-world cases predate generative AI, underscoring that unlawful behaviors may arise even with simpler technologies – and may reemerge, potentially amplified, with more autonomous and knowledgeable agents.
Table 1: Examples of Unlawful or Harmful Behavior in Multi-Agent AI Systems, selected from HammondMultiAgentRisksAdvanced2025 and deWittOpenChallengesMultiAgent2025.
@@ 3.3 Mapping Risks in Multi-Agent AI Systems
Here, I provide a compact taxonomy of these risks. The list is not exhaustive: its purpose is to offer readers unfamiliar with AI systems and agents a first overview of the main plausible sources of harm, while pointing to more detailed surveys for technical depth (HammondMultiAgentRisksAdvanced2025; deWittOpenChallengesMultiAgent2025; BengioSuperintelligentAgentsPose2025). The risks are diverse and heterogeneous, spanning development processes, decision-making dynamics, and institutional responses. They cut across disciplinary boundaries, underscoring the importance of transdisciplinary integration to meaningfully anticipate and mitigate them. [^18] For completeness, in Section A of the Appendix I also elaborate on the benefits of this increasingly plausible socio-technical horizon, to provide a more balanced perspective on this transformative transition, underscoring how Multi-agent AI systems should not be seen exclusively as potential generators of harm.
@@
Social Learning Theory itself has long explained how deviant behavior in humans often stems from social interaction. Peer groups, family, and colleagues can promote either conformity or deviance, depending on the reinforcement environment (WarrInfluenceDelinquentPeers1991; SimonsLearningBeBad2011; AkersSocialLearningSocial2017). The same logic may apply to machines interacting autonomously with each other: agents not originally designed for harm may, through interaction, adopt negative behaviors via mechanisms such as imitation and reinforcement (XieWhosMoleModeling2025).
@@
Additionally, scholarship on social networks shows how interactions accelerate the diffusion of ideas and behaviors, positive or negative alike (Bakshyrolesocialnetworks2012; KimSocialnetworktargeting2015; CinelliCOVID19socialmedia2020). Just as pathogens spread faster in highly connected populations (GlassSocialcontactnetworks2008; ClipmanDeeplearningsocial2022), harmful behaviors could proliferate more quickly in tightly coupled multi-agent systems than in isolated ones.
@@
From a monitoring and intervention perspective, identifying the causes of harmful behavior within such systems becomes substantially more difficult. Understanding which agent initiated a harmful act, how it spread, and through which pathways requires robust methods of causal inference. Yet, causal discovery in networked systems is notoriously complex, especially under interference and feedback conditions (VanderWeeleSocialNetworksCausal2013; SussmanElementsestimationtheory2017; MaCausalInferenceNetworked2021; ClipmanDeeplearningsocial2022). As such, interacting agent systems risk becoming “two-layered black boxes”: one opaque layer within each agent, and another arising from the system of interactions itself.
@@
Connected to the previous point, as the complexity of multi-agent systems increases, so too does the challenge of interpreting, auditing, and ultimately controlling their behavior (BansalEmergentComplexityMultiAgent2018; GrupenConceptbasedUnderstandingEmergent2022). Inter-agent interactions can create feedback loops, conditional dependencies, and non-linear effects that obscure the logic of any given action or decision. The result is a system that may behave in ways that are technically functional but epistemically opaque. This opacity not only complicates efforts to ensure accountability but also undermines user trust, particularly in domains where transparency is a legal or ethical requirement. In this regard, a system of interacting AI agents may become more than the sum of its parts: it may become a fundamentally alien system from the standpoint of human interpretability.
@@
Legal and regulatory challenges would also emerge in multi-agent AI systems. Existing frameworks for responsibility and liability are poorly suited for multi-agent dynamics (CerkaLiabilitydamagescaused2015; TurnerRobotRulesRegulating2018; PricePotentialLiabilityPhysicians2019). As more actors – human or non-human – become entangled in decision-making chains, assigning accountability for harmful outcomes becomes increasingly ambiguous. In parallel, ensuring smooth, effective governance also represents a challenge in this context (DignumResponsibleAIAutonomous2025). The governance dimension entails virtually every aspect concerned with the engineering and deployment of multi-agent systems: how can we design sustainable and effective oversight procedures? Which institutions, in a highly globalized world and in borderless digital domains, will be responsible for monitoring these systems? What role should private companies play in this process? These are some of the key questions that demand attention, inherently linking regulation and governance together.
@@
Multi-agent interaction may be vulnerable to adversarial exploitation. In a future where even critical infrastructures are governed by interacting AI agents, malicious actors could induce large-scale disruption by targeting systemic vulnerabilities. This risk mirrors the logic of cascading failures, studied in relation to power grids and financial systems (ZhaoSpatiotemporalpropagationcascading2016; YangSmallvulnerablesets2017; SchaferDynamicallyinducedcascading2018; BaqaeeCascadingFailuresProduction2018), but differs in that interactive agents may possess adaptive capabilities, making their behavior less predictable and more difficult to control.
@@
Coordination failures, unintended competition, or outright conflict may emerge when agents operate with overlapping but unaligned goals, or when resource constraints lead to strategic divergence (HammondMultiAgentRisksAdvanced2025; PanWhyMultiagentSystems2025). In such contexts, agents may begin to exhibit adversarial behaviors, competing for access to data, processing resources, or strategic positioning. These failures can degrade performance and, in some cases, produce socially harmful outcomes. The risk of such breakdowns increases in systems lacking explicit coordination protocols or oversight mechanisms, especially when deployed in open or decentralized environments.
@@
Finally, the performance of multi-agent systems may not scale linearly with the number of agents involved. As agent populations grow, the complexity of the system’s internal dynamics may increase exponentially, leading to emergent forms of instability (MaEfficientscalablereinforcement2024). These can manifest as oscillations, feedback-driven runaway behaviors, or systemic fragility, dynamics that are difficult to predict or preempt. This is especially problematic in infrastructure systems or critical services, where failures can propagate rapidly and non-locally. In this light, the move toward interacting agent populations must be accompanied by a serious effort to model and anticipate second-order effects that arise specifically at scale.